OpenStudy (anonymous):

Solve the system algebraically. 4x - 5y = 3 x = 3 - y What is the value of x?

OpenStudy (anonymous):

Lets break it up. Taking just the first equality: 4x - 5y = 3x x - 5y = 0 x = 5y y = x/5 Then the second equality tell us that 3x = 3-y Using y = x/5 that we got before here gives 3x = 3 - (x/5) Can you get take from there?